# SAE-V: Interpreting Multimodal Models for Enhanced Alignment

### v1

@@ 4.2 Experiment Results
Figure 7: The performance of SAE-V-based data filter method SAE-V-based data filter method significantly outperforms the random selection baseline. Specifically, the cosine similarity filter demonstrated the most stable performance, consistently surpassing the random filter across all data percentages and achieving 108% of the full dataset’s performance with only 20% of the data. The co-occurrence filter peaked at 50% of the data, reaching a score of 108.17, (115% of the full dataset’s performance). As a more straightforward utilization of SAE-V model, the L0 filter also generally outperforms the random selection baseline.
@@ 4.4 Ablation Study
Figure 9: The performance of SAE-V-based data filter method on Chameleon-7B We replicate SAE-V-based data filter on Chameleon-7B model and Align-Anything dataset. The result shows that on Chameleon-7B model, SAE-V-based data filter also achieved significant performance gain compared to the random filter. Specifically, both cosine similarity and co-occurrence filters perform better than random filters at almost every data percentage, and the cosine similarity filter achieved a score of 52.13 (120% of the full dataset’s performance) with 70% data.
@@
To prove that SAE-V-based data filter method could generalize to distinct model architectures, we replicate SAE-V-based data filter on Chameleon-7B model and Align-Anything dataset. The result (shown in Figure 9) demonstrates that although Chameleon-7B performs worse than LLaVA-NeXT-7B on LLaVA-Bench, the SAE-V-based filter method still shows its effectiveness. When using a smaller data proportion, the performance is strongly correlated with the data quantity, and thus the differences between methods are minimal. However, with a larger data proportion, the SAE-V-based filter method significantly surpasses the random filter, achieving a peak of 52.13 (120% of the full dataset’s performance) with 70% of the data. The largest performance gap is observed in the 50-70% data range, while the differences converge again as the data proportion approaches 100%. This proves that SAE-V-based data filter is effective on architectures other than CLIP-based MLLM, and shows its potential to generalize across a wide range of models.
Figure 10: The performance of SAE-V-based data filter method on RLAIF-V dataset. We performed an ablation study with the RLAIF-V dataset and Chameleon-7B model to prove that SAE-V-based data filter method can generalize across datasets. Both cosine similarity and co-occurrence filters generally outperform the random filter on the RLAIF-V dataset. Specifically, the cosine similarity filter achieves the highest score of 54.23 (125% of the full dataset’s performance) with only 10% data, demonstrating its supreme efficiency.
@@
We also performed an ablation study on the datasets to be filtered. Since the LLaVA-NeXT-7B model is highly capable, most datasets fail to further enhance its multimodal abilities. Therefore, we selected the RLAIF-V dataset and the relatively weaker Chameleon-7B model for the dataset ablation study. The results (shown in Figure 10) further confirm that SAE-V-based data filter is working across different datasets. Moreover, on RLAIF-V, the cosine similarity filter could achieve a score of 54.23 (125% of the full dataset’s performance) by using only 10% of the data, demonstrating exceptional efficiency.
@@
To validate the effectiveness of our SAE-V-based data filtering method, we conducted an ablation study comparing it with other similar data filtering approaches. Since there are currently no widely recognized data filtering methods specifically designed for multimodal data, we adapted the IFD metric (Li et al., 2023b) method to the multimodal setting. The result (shown in Figure 11) suggests that our data filter method achieves a performance comparable to the IFD metric. However, considering that the IFD metric needs to train an additional cherry model, our SAE-V-based data filter could directly fit various datasets, demonstrating greater generalizability and efficiency.
@@ 5 Related Work
  • CLIP-based MLLMs: These models encode images with CLIP (Radford et al., 2021) and use MLP to project visual features. Examples include LLaVA (Liu et al., 2024) series and NExT-GPT (Wu et al., 2023).
@@
  • Q-Former-based MLLMs: These models use a structure similar to Q-Former (Li et al., 2023a) to extract visual representations, represented by Qwen-VL (Bai et al., 2023) and MiniGPT-4 (Zhu et al., 2024).
@@
Sparse Autoencoder (SAE) aims to learn sparse and interpretable features from polysemantic model representations (Yun et al., 2021; Bricken et al., 2023; Sharkey et al., 2022; Peigné, 2023; Elhage et al., 2022). By introducing sparsity constraints, the activation values in the hidden layers of SAE are mostly zero, allowing SAE to encode polysemantic features in LLM to monosemantic ones.
@@
Data filtering ensures that only relevant high-quality data are used during the alignment of LLM or MLLM s, thus reducing the quantity of data while achieving greater performance (Zhou et al., 2023; Chen et al., 2023; Du et al., 2023; Li et al., 2023c, b; Tu et al., 2024). For example, LIMA (Zhou et al., 2023), ALPAGASUS (Chen et al., 2023), and IFD (Li et al., 2023b) use human annotation, API annotation and train a new model for annotation to score data separately. Our method, SAE-V-based data filter, provides a self-guided and interpretable metric to evaluate the similarity of multimodal data, which indicates their qualities. The method is stable and efficient for models of different architectures.
@@ A.1 Hyperparameter of SAE and SAE-V Models Training
The hyperparameters of the training SAE and SAE-V are shown in Table 2. The differences in training parameters arise because the LLaVA-NeXT-7B model requires more GPU memory to handle vision input, so fewer batches can be cached. For the SAE and SAE-V parameters, we set different hook layers and context sizes based on the distinct architectures of the two models. We also experimented with different feature numbers on both models, but found that only around 30,000 features are actually activated during training. All training runs were conducted until convergence. All SAE and SAE-V training is performed on 8×A800 GPUs. We ensured that the variations in the parameters did not affect the experiment results.
@@ B.1 Algorithms in alignment experiment
     $F_{i}\leftarrow\{\bm{f}_{k}\text{:}\|\bm{f}_{k}Z_{i}^{\top}\|_{\infty}>\delta\}$
@@
        $\mathcal{A}_{k}\leftarrow\mathcal{A}_{k}\cup\{\bm{h}_{j}\text{:}\ \bm{h}_{j}\in H_{i},\bm{z}_{j}=\bm{e}_{j}Z_{i},\bm{z}_{j}\bm{f}_{k}^{\top}>\delta\}$
@@ C.1 Algorithm
     $\mathcal{A}_{j}\leftarrow\mathcal{A}_{j}\cup\{\bm{h}_{i}\text{:}\ \bm{h}_{i}\in H,\bm{z}_{i}=\bm{e}_{i}Z,\bm{z}_{i}\bm{f}_{j}^{\top}>\delta\}$
@@
     $\mathcal{A}_{j}\leftarrow\mathcal{A}_{j}\cup\{\bm{h}_{i}\text{:}\ \bm{h}_{i}\in H,\bm{z}_{i}=\bm{e}_{i}Z,\bm{z}_{i}\bm{f}_{j}^{\top}>\delta\}$

### v2

@@ 4.3 Relationship between MLLM Capability and SAE-V Features
Method 0% 10% 20% 30% 40% 50% 60% 70% 80% 90% 100%
SAE-V 94.2 98.2 106.8 114.9 114.5 114.8 112.9 112.3 111.0 109.5 98.5
Random 94.2 96.5 97.0 98.3 95.3 93.7 96.5 98.8 98.2 96.8 98.5
Table 2: The performance of SAE-V-based data filter method on 1/20 Align-Anything dataset The performance of SAE-V-based data filter method on 1/20 Align-Anything dataset closely resembles that in Figure 9, confirming that alignment metrics on a small validation set resemble the distribution on the complete dataset, enabling efficient hyperparameter selection.
Method Performance
0% 20% 40% 60% 80% 100%
SAE-V 104.6 105.8 116.7 112.3 112.0 111.2
Random 104.6 105.3 105.8 107.2 110.4 111.2
Table 3: The performance of SAE-V-based data filter method on LLaVA-NeXT-Vicuna-13B We replicate SAE-V-based data filter on LLaVA-NeXT-Vicuna-13B and Align-Anything dataset. The result shows that SAE-V-based data filter outperforms the random selection baseline, and reached the highest performance of 116.67 with 40% data. This further indicates that SAE-V-based data filter could work regardless of model architecture and sizes.
@@ 4.4 Ablation Study
To prove that SAE-V-based data filter method could generalize to distinct model architectures, we replicate SAE-V-based data filter on Chameleon-7B and LLaVA-NeXT-Vicuna-13B models with the Align-Anything dataset. The result (shown in Figure 8 (a) and Table 3) demonstrates that SAE-V-based filter method still shows its effectiveness regardless of model architecture, sizes, and image encoding methods.
When using a smaller data proportion, the performance is strongly correlated with the data quantity, and thus, the differences between methods are minimal. However, with a larger data proportion, the SAE-V-based filter method significantly surpasses the random filter, achieving a peak of 52.13 (120% of the performance on full dataset) with 70% data on Chameleon-7B and 116.67 (104% of the performance on full dataset) with 40% data onLLaVA-NeXT-Vicuna-13B. The largest performance gap is observed in the 40-80% data range, while the differences converge again as the data proportion approaches 100%. This proves that SAE-V-based data filter is effective on architectures other than CLIP-based MLLM, and shows its potential to generalize across a wide range of models.
@@
We also performed an ablation study on the datasets to be filtered. Since the LLaVA-NeXT-7B model is highly capable, most datasets fail to further enhance its multimodal abilities. Therefore, we selected the RLAIF-V (Yu et al., 2024) and MMInstruct (Liu et al., 2024b) datasets with the relatively weaker Chameleon-7B model for the dataset ablation study. The results (shown in Figure 8 (b) and Figure 8 (c)) further confirm that SAE-V-based data filter is working across different datasets. Moreover, on RLAIF-V, the cosine similarity filter could achieve a score of 54.23 (125% of the full dataset’s performance) by using only 10% of the data, and on MMInstruct, the cosine similarity filter could achieve a score of 61.2 (117% of the full dataset’s performance) with 60% data, demonstrating exceptional efficiency.
Ablation on Dataset Sizes
On real deployment scenarios, due to the large size of datasets, it is often difficult to use the SAE-V-based data filter method to search for filtering ratio parameters on the complete dataset. Therefore, we reproduced the SAE-V-based data filter method on a 1/20 subset of the Align-Anything dataset to demonstrate that in practical deployment situations, it is feasible to search for filter ratios on small-scale dataset subsets for application to the complete dataset. The results (shown in Table 2) indicate that the performance of SAE-V-based data filter on the small-scale dataset exhibits extremely similar trends with respect to filtered data ratios as those observed on the complete dataset (shown in Figure 9), which demonstrates that searching for filtered data ratios through small-scale dataset subsets is feasible.
@@
To validate the effectiveness of our SAE-V-based data filtering method, we conducted an ablation study comparing it with other similar data filtering approaches. Since there are currently no widely recognized data filtering methods specifically designed for multimodal data, we adapted the IFD metric (Li et al., 2024a) method to the multimodal setting. The result (shown in Figure 11) suggests that our data filter method achieves a performance comparable to the IFD metric. However, considering that the IFD metric needs to train an additional cherry model, our SAE-V-based data filter could directly fit various datasets, demonstrating greater generalizability and efficiency.
@@ 5 Related Work
  • CLIP-based MLLMs: These models encode images with CLIP (Radford et al., 2021) and use MLP to project visual features. Examples include LLaVA (Liu et al., 2024a) series and NExT-GPT (Wu et al., 2024b).
@@
  • Q-Former-based MLLMs: These models use a structure similar to Q-Former (Li et al., 2023) to extract visual representations, represented by Qwen-VL (Bai et al., 2023) and MiniGPT-4 (Zhu et al., 2024).
@@
Sparse Autoencoder (SAE) aims to learn sparse and interpretable features from polysemantic model representations (Yun et al., 2021; Bricken et al., 2023; Sharkey et al., 2022; Peigné, 2023; Elhage et al., 2022). By introducing sparsity constraints, the activation values in the hidden layers of SAE are mostly zero, allowing SAE to encode polysemantic features in LLM to monosemantic ones. Zhang et al. (2024) firstly attempted to apply SAE to analyze the open-semantic features of MLLMs.
@@
Data filtering ensures that only relevant high-quality data are used during the alignment of LLMs or MLLMs, thus reducing the quantity of data while achieving greater performance (Zhou et al., 2023; Chen et al., 2024; Du et al., 2023; Li et al., 2024b, a; Tu et al., 2025). For example, LIMA (Zhou et al., 2023), ALPAGASUS (Chen et al., 2024), and IFD (Li et al., 2024a) use human annotation, API annotation and train a new model for annotation to score data separately. Our method, SAE-V-based data filter, provides a self-guided and interpretable metric to evaluate the similarity of multimodal data, which indicates their qualities. The method is stable and efficient for models of different architectures.
@@ 6 Conclusion
Acknowledgement
This work is sponsored by National Natural Science Foundation of China (62376013, 623B2003).
@@ A.1 Hyperparameter of SAE and SAE-V Models Training
The hyperparameters of the training SAE and SAE-V are shown in Table 4. The differences in training parameters arise because the LLaVA-NeXT-7B model requires more GPU memory to handle vision input, so fewer batches can be cached. For the SAE and SAE-V parameters, we set different hook layers and context sizes based on the distinct architectures of the two models. We also experimented with different feature numbers on both models, but found that only around 30,000 features are actually activated during training. All training runs were conducted until convergence. All SAE and SAE-V training is performed on 8×A800 GPUs and each training typically takes around $21$ hours. We ensured that the variations in the parameters did not affect the experiment results.
@@ B.1 Algorithms in alignment experiment
     $F_{i}\leftarrow\{\bm{f}_{k}\text{: }\exists\ z_{j}\in Z_{i},\ \mathrm{s.t.}\ z_{j}\ \text{activates }\bm{f}_{k}\}$
@@
        $\mathcal{A}_{k}\leftarrow\mathcal{A}_{k}\cup\{\bm{h}_{j}\text{:}\ \bm{h}_{j}\in H_{i},\bm{z}_{j}=\bm{e}_{j}Z_{i},\bm{z}_{jk}>\delta\}$
@@ C.1 Algorithm
     $\mathcal{A}_{j}\leftarrow\mathcal{A}_{j}\cup\{\bm{h}_{i}\text{:}\ \bm{h}_{i}\in H,\bm{z}_{i}=\bm{e}_{i}Z,\bm{z}_{ij}>\delta\}$
@@
     $\mathcal{A}_{j}\leftarrow\mathcal{A}_{j}\cup\{\bm{h}_{i}\text{:}\ \bm{h}_{i}\in H,\bm{z}_{i}=\bm{e}_{i}Z,\bm{z}_{ij}>\delta\}$
@@ C.2 Case Analysis
C.3 Quantative Analysis
Masking (%) 0 25 50 75
Text Accuracy (%) 80.2 77.7 66.5 53.3
Image Accuracy (%) 80.2 78.8 78.4 70.7
Table 7: Applying SAE-V on VQA tasks We apply SAE-V-based patch filtering experiment on the text and image of VQA task separately, using A-OKVQA validation set with LLaVA-NeXT-7B model. The experiment shows that image information demonstrates a lower compression rate (more redundancy) than text. Moreover, text masking shows a roughly linear relation of accuracy and masked token, while image filter maintains performance until 50%, with a more significant drop only appearing at 75%, suggesting that the information of text is more evenly distributed compared to image.
We performed additional experiments on SAE-V-based patch filtering. To be specific, we use SAE-V-based patch filter on the text and image part of each VQA question separately, and test the accuracy when part of the information is masked. The results show that image information demonstrates a lower compression rate (more redundancy) than text. Moreover, text masking shows a roughly linear relation of accuracy and masked token, while image filter maintains performance until 50%, with a more significant drop only appearing at 75%, suggesting that the information of text is more evenly distributed compared to image.